OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y less than or greater to x - 5 y greater than or equal to -x - 4 (-5, 2) (5, -2) (-5, -2) (5, 2)

OpenStudy (anonymous):

please help me

OpenStudy (jdoe0001):

well, test the points against the inequalities let's check the 1st one y <= x-5 (-5, 2) (2) <= (-5)-5 2 <= 0 # is that true, well, 2 is really bigger than 0, so no dice there now test the 2nd point

OpenStudy (anonymous):

oh i see it will be (-5,2) thats the answer right?

OpenStudy (jdoe0001):

after replacing the "y" and "x" with their corresponding values in the ordered set given (-5, 2) ..( ...) x, y the inequality must be true for both cases testing the 1st point, it gave \(\bf 2 \le 0 \) which is not true, thus it's not the solution

OpenStudy (anonymous):

im confuse so what the answer

OpenStudy (jdoe0001):

well, test B see if it works, if it doesn't, test C

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i think answer is (-5,2)

OpenStudy (jdoe0001):

we test it that one, we got as answer \(\bf 2 \le 0\) now, is 2 really smaller than 0?

OpenStudy (jdoe0001):

or equal for that matter

OpenStudy (anonymous):

soo (-5,-2)

OpenStudy (jdoe0001):

let's test B 1st case y <= x - 5 point is ( 5, -2) x y what does it give you?

OpenStudy (anonymous):

(5,-2)

OpenStudy (anonymous):

well can u just give me the answer

OpenStudy (jdoe0001):

anything stopping you from testing it?